A person consumes about 2900Cal a day. 1 Cal = 4186 J , and 1 kWh = 860 Cal .

A) What is the energy in joules? Q=? J

B) What is this energy in kilowatt-hours? Q=? kWh

C) If a power company charges about 14
Answer #1

A)energy in jouls=2900*4186=12139400 J

B)energy in kWH=2900/860=3.3721 KWh

C) energy consumption per day =3.3721 kWh

for 1KWh charge is 14 $

for 3.3721KWh it is 3.3721 *14=47.21 $
Answer #2

Energy=2900 Cal=2900 *4186=Kwh12139400 J
B)E=2900 cal =2900 *(1/860)=3.37 Kwh
c)The cost per day =3.37*14 =47.29 $